# Pairwise vs Higher-order interactions: Can we identify the interaction type in coupled oscillators from time series?

### v1

@@ Abstract
Rhythmic phenomena, which are ubiquitous in biological systems, are typically modelled as systems of coupled limit cycle oscillators. Recently, there has been an increased interest in understanding the impact of higher-order interactions on the population dynamics of coupled oscillators. Meanwhile, estimating a mathematical model from experimental data is a vital step in understanding the dynamics of real-world complex systems. In coupled oscillator systems, identifying the type of interaction (e.g. pairwise or three-body) of a network is challenging, because different interactions can induce similar dynamical states and bifurcations. In this study, we have developed a method based on the adaptive LASSO (Least Absolute Shrinkage and Selection Operator) to infer the interactions between the oscillators from time series data. The proposed method can successfully classify the type of interaction and infer the probabilities of the existence of pairwise and three-body couplings. Through systematic analysis of synthetic datasets, we have demonstrated that our method outperforms two baseline methods, LASSO and OLS (Ordinary Least Squares), in accurately inferring the topology and strength of couplings between oscillators. Finally, we demonstrate the effectiveness of the proposed method by applying it to the synthetic data of 100 oscillators. These results imply that the proposed method is promising for identifying interactions from rhythmic activities in real-world systems.
@@ Introduction
A complex system can be modelled as a network of interacting nodes or vertices (i.e. elements) connected by links or edges (i.e. individual interaction or couplings). However, in many systems, such as social, ecological, and biological systems, the effects of interaction between groups of three or more elements (higher-order interaction) cannot be ignored. Recently, there has been a significant amount of research on the effects of higher-order interaction between the elements on the network dynamics [17, 18]. In coupled oscillator systems, it has been reported that higher-order interaction can emerge from pairwise interaction, e.g., via the nonlinearity of the coupling [19, 20] and the dependence of the amplitude on the phase [21]. In addition, higher-order interactions can arise spontaneously due to frequency resonance of oscillators [22]. There is also extensive research on the effects of higher-order interactions on dynamics [17]. In coupled phase oscillators, three-body interactions can also induce synchronous states, asynchronous states, and cluster states [23, 24, 25], similar to pairwise interactions [26]. In addition, three-body interactions exhibit explosive synchronization, associated with the bistabilisation of synchronous states [20, 23], in a manner analogous to pairwise interactions [27, 28]. Consequently, both pairwise and three-body interactions can induce similar dynamics and bifurcations in coupled oscillator systems.
@@
Estimating a mathematical model from experimental data is an essential step in understanding the dynamics of complex systems in the real world. In the context of coupled oscillator systems, previous studies have developed methods for identifying phase models from data. These methodologies include estimating the phase response curves [29, 30, 31], phase coupling functions [32, 33, 34, 35] (see also review [36, 37]), and asymptotic phases [38, 39, 40]. While several studies have proposed methods for estimating the interaction of coupled oscillator systems [41, 42, 43, 44], these are based on the assumption of pairwise interactions. Recently, new methods for estimating the higher-order interaction in complex dynamical systems have been proposed [45, 46]. However, it remains challenging to identify the type of interaction, i.e. whether the elements are coupled via two-body pairwise interaction or three-body or higher order group interaction, only from data. In coupled oscillator systems, it is not possible to determine the type of interaction solely from their qualitative collective behaviour, such as synchronisation or clustering or from their bifurcations. It is therefore essential to develop a method for identifying the type of interaction and inferring the topology and strength of the couplings between the oscillators.
@@
In this study, we address the following question: "Is it possible to identify the interaction type of coupled oscillators from the data?" To this end, we develop a method based on the adaptive LASSO for inferring the topology and strength of the couplings (i.e. individual interaction) between the oscillators from time series. We then systematically evaluate the performance of the proposed method by analysing synthetic datasets with different interaction types, coupling probabilities and observation durations, and compare the performance with baseline methods. Finally, we demonstrate the applicability of the proposed method by inferring the interaction of a system of 100 oscillators.
Result
@@
Figure 1: Synchronization dynamics of weakly coupled oscillators in the presence of white noise (Eq. 1). Left: Time series of the Kuramoto order parameter $r$ in a weakly synchronized state for three types of interactions; A. Pairwise interaction only, B. Three-body interaction only, and C. Mixture interactions, where the system has both pairwise and three-body couplings. The average values of the order parameter were $r=0.33,0.25$, and $0.32$ for the pairwise, three-body, and mixture interactions, respectively. Right: Snapshots of the phase distribution of the oscillators when they are synchronized at similar levels ($r=0.67,0.61$ and $0.67$ for the pairwise, three-body, and mixture interaction, respectively). The weights ($K_{2}$, $K_{3}$) and the coupling probabilities ($q^{(2)}$, $q^{(3)}$) were set as follows: A. $K_{2}=0.1$, $q^{(2)}=0.1$and $q^{(3)}=0.0$, B. $K_{3}=0.3$, $q^{(2)}=0.0$, and $q^{(3)}=0.1$, and C. $K_{2}=0.1$, $K_{3}=0.3$, $q^{(2)}=0.05$ and $q^{(3)}=0.05$.
@@
The adaptive LASSO method for inferring the interaction network from time series
@@
In this paper, we develop a method based on the adaptive LASSO [47] for inferring the interaction parameters from time series. The model parameters, including the natural frequency ($\omega_{i}$) and the coupling parameters (${\bf W}^{(2)}_{i}$ and ${\bf W^{(3)}_{i}}$), are determined by fitting the dynamical model (1) for an oscillator $i$. All parameters are determined by solving this problem repeatedly for all oscillators ($i=1,2,\cdots,N$). First, a preliminary estimate of the parameters of the $i$-th oscillator ($\omega_{i}$, ${\bf W}^{(2)}_{i}$, and ${\bf W}^{(3)}_{i}$) is obtained by the OLS (Ordinary Least Squares) method [45]. These parameters are determined through minimizing the squared error given by
@@
where $h$ is the sampling interval. The estimate is then refined through minimizing the cost function given by
$E_{\rm AL}=E_{\rm OLS}+\alpha\left\{\sum_{j:j\neq i}\left|c^{(2)}_{ij}W^{(2)}_{ij}\right|+\sum_{j,l:l>j,j\neq i,l\neq i}\left|c^{(3)}_{ijl}W^{(3)}_{ijl}\right|\right\},$ (5)
where $|x|$ represents the absolute value of a real number $x$, and $c^{(2)}_{ij}$ and $c^{(3)}_{ijl}$ are hyperparameters, which are set as follows: $c^{(2)}_{ij}=1/\tilde{W}^{(2)}_{ij}$ and $c^{(3)}_{ijl}=1/\tilde{W}^{(3)}_{ijl}$, where $\tilde{W}^{(2)}_{ij}$ and $\tilde{W}^{(3)}_{ijl}$ are the preliminary estimates obtained by OLS (Eq. 4). A hyperparameter $\alpha$ is determined by minimizing Bayesian Information Criterion (BIC), which is performed by using LassoLarsIC function from scikit-learn [48]. This adaptive LASSO method is capable of distinguishing whether or not there is a coupling between a pair or a triple of the nodes. If a coupling is identified ($W^{(2)}_{ij}\neq 0$ or $W^{(3)}_{ijl}\neq 0$), this method is also able to infer the coupling strengths.
Figure 2 compares the inferred couplings from the proposed method with those from two baseline methods, LASSO (Least Absolute Shrinkage and Selection Operator) and OLS (Ordinary Least Squares) with a statistical test (see Method for details). The proposed method was able to infer the topology of the pairwise and three-body couplings with complete accuracy (Fig. 2, Proposed), with no false positives and no false negatives. OLS produced a considerable number of false positives for both pairwise and three-body couplings (Fig. 2, OLS). When the true interaction was pairwise only, the number of false positives was 16 (12 %) and 23 (3.5 %) for pairwise and three-body couplings, respectively. While LASSO gave a lower number of false positives than OLS, it gave the false positives when the true interaction was pairwise (Fig. 2, LASSO). When the true interaction was pairwise only, the number of false positives was 0 (0 %) and 12 (1.8 %) for pairwise and three-body couplings, respectively. Furthermore, the width of the arrow indicates that LASSO underestimates the coupling strength. These results suggest that the proposed method performs better than OLS and LASSO in inferring the couplings from the time series. In the next subsection, we systematically compare the inference performance of the proposed method with that of the baseline methods.
@@
Table 1: Accuracy of inferring the interaction type (pairwise, three-body, or mixture) from time series. ’Proposed’ represents the accuracy of the proposed method. ’LASSO’ represents the accuracy of the LASSO method. ’OLS’ and ’OLS (Naive)’ represent the accuracy of OLS with a statistical test and OLS with naive thresholding, respectively (see Methods for detail of LASSO and OLS methods). Bold letters indicate the most accurate method. The synthetic data were generated from 12 oscillators.
Interaction type Proposed LASSO OLS OLS (Naive)
Pairwise 100 % 100 % 100 % 100 %
Three-body 100 % 100 % 100 % 80 %
Mixture 100 % 100 % 85 % 85 %
Table 2: Inference of the probabilities of the existence of pairwise and three-body couplings from phase time series. ’Proposed’ represents the result of the proposed method. ’LASSO’ represents the result of the LASSO method. ’OLS’ and ’OLS (Naive)’ represent the result of OLS with a statistical test and OLS with naive thresholding, respectively (see Methods for detail of LASSO and OLS methods). Bold letters indicate the most accurate method. The synthetic data were generated from 12 oscillators.
True Proposed LASSO OLS OLS (Naive)
Pairwise $q^{(2)}$ 10.0 % 10.4 % 7.7 % 16.1 % 23.1 %
$q^{(3)}$ 0.0 % 0.1 % 0.7 % 0.6 % 9.2 %
Three-body $q^{(2)}$ 0.0 % 0.8 % 0.0 % 0.1 % 5.1 %
$q^{(3)}$ 10.0 % 10.1 % 10.8 % 8.3 % 13.7 %
Mixture $q^{(2)}$ 4.9 % 5.6 % 4.3 % 6.1 % 12.2 %
$q^{(3)}$ 4.9 % 4.9 % 5.4 % 4.2 % 10.2 %
@@
Table 1 shows the accuracy of inferring the interaction type from the time series, which were generated from 20 populations of each pairwise, three-body, and mixture interaction. All the methods achieve an accuracy of over 80 % in identifying the interaction type. In particular, the proposed method and LASSO achieve perfect identification of the interaction type, i.e. 100 % accuracy. Table 2 shows the inferred probabilities of the existence of pairwise and three-body couplings ($\hat{q}^{(2)}$ and $\hat{q}^{(3)}$) obtained from the same data as in Table 1. In the case of synthetic data with pairwise interaction, the proposed method demonstrates the highest accuracy in inferring the coupling probability. In contrast, LASSO underestimated and OLS overestimated the probability of pairwise coupling. In addition, both LASSO and OLS exhibited an overestimation of the probability of three-body coupling by 0.6-0.7 %, and the naive OLS exhibited an overestimation of more than 9.0 %. In the case of synthetic data with three-body interaction, LASSO was the most accurate in inferring the probability of the pairwise couplings, while the proposed method was the most accurate in inferring that of the three-body couplings. In contrast, Naive OLS overestimated the probability of pairwise (5.1 %) and three-body (3.7 %) couplings, and OLS underestimated the probability of the three-body couplings. In the case of synthetic data with mixture interaction, LASSO achieved the most accurate inference for the probability of the pairwise couplings, while the proposed method achieved the most accurate inference for that of the three-body couplings. OLS overestimates the probability of the pairwise couplings, and it underestimates that of the three-body couplings. Naive OLS overestimated the probability of pairwise (7.3 %) and three-body (5.3 %) couplings. Overall, the naive OLS method cannot infer the coupling probabilities in the present situation. In contrast, a recent study [45] has shown that OLS method can infer the higher-order interactions from time series accurately. The discrepancy between these findings can be attributed to the difference in the situation being considered. Previous research has focused on scenarios without noise, while this study considers a system of noisy coupled oscillators. Tables 2 indicates that sophisticated methods, including OLS with a statistical test and LASSO [49, 47] (see Method for details), are essential for distinguishing between weak couplings and no couplings in noisy situations. In the following, the proposed method is compared with two sophisticated methods, i.e. OLS with statistical test and LASSO.
@@
where $N_{\rm TP}$, $N_{\rm TN}$, $N_{\rm FP}$, and $N_{\rm FN}$ are the numbers of true positive, true negative, false positive, and false negative of the pairwise and three-body couplings, respectively. MCC values range from -1 to 1. If all the couplings are correctly inferred, the MCC value is 1. If the couplings are randomly inferred with the true coupling probability, the expected value of MCC is 0. Figure 3 shows the dependence of the MCC on the observation duration for the three interaction types (pairwise, three-body, and mixture) with three coupling probabilities: $q=0.05,0.10$, and $0.15$. As the coupling probability increases, a longer observation period is required for an accurate inference. For example, 50 cycles are sufficient to accurately ($MCC>0.8$) infer the couplings for the synthetic data of a sparse interaction ($p=0.05$). In contrast, more than 100 cycles are necessary to infer the couplings for the synthetic data of a dense interaction ($p=0.15$). Overall, when the observation period is sufficiently long, the proposed method demonstrates superior performance in terms of accuracy compared to LASSO and OLS. Furthermore, the proposed method outperforms LASSO and OLS in inferring the couplings across a broad range of coupling strengths, $K_{2}$ and $K_{3}$ (see Supplementary Fig. 1).
@@
Figure 3: Effect of observation time on inference accuracy. Performance of coupling inference (MCC) was calculated for the three types of interaction: A. Pairwise ($q^{(2)}=q$, $q^{(3)}=0$), B. Three-body ($q^{(2)}=0$, $q^{(3)}=q$), and C. Mixture of pairwise and three-body interactions ($q^{(2)}=q/2$, $q^{(3)}=q/2$). The parameter of the coupling probability $q$ was set to 0.05, 0.10 and 0.15, on the left, centre and right panel, respectively.
Table 3: Accuracy of coupling inference measured by MCC. Bold letters indicate the best method.
Interaction type Proposed LASSO OLS
Pairwise 0.92 0.59 0.60
Three-body 0.85 0.83 0.81
Mixture 0.85 0.83 0.78
Table 4: Accuracy of coupling strength inference measured by mean squared error. Bold letters indicate the best method.
Interaction type Proposed LASSO OLS
Pairwise 0.012 0.056 0.239
Three-body 0.019 0.032 0.138
Mixture 0.013 0.028 0.162
@@
Finally, we evaluate the inference performance obtained from the coupled oscillators with highly heterogeneous coupling strength. Such interactions are observed in neural circuits in the brain [51, 50]. We simulate a population of oscillators (1) whose coupling strength is given by the Erdös-Rényi random graph with a log-normal distribution of the strength. Figure 4 shows the inference result of the coupling strength by the three methods (our proposed method, LASSO, and OLS). True negatives are plotted at the origin, false negatives are plotted on the positive part of the $x$-axis, false positives are plotted on the positive part of the $y$-axis, and true positives are plotted in the first quadrant. Again, the performance of coupling inference is evaluated using MCC (Table 3). The MCC values of the proposed method were higher than those of LASSO and OLS, indicating that the proposed method outperforms LASSO and OLS in terms of coupling inference. Furthermore, we examine the performance in inferring the coupling strengths. The coupling strengths inferred by OLS are distributed above the diagonal line, indicating that OLS overestimates the strength. Conversely, the coupling strengths inferred by LASSO are distributed below the diagonal line, indicating that LASSO underestimates the strength. The performance in inferring the strength is evaluated by calculating the mean squared error: $\epsilon^{2}=\frac{1}{M}\sum_{C}(w_{C}-\hat{w}_{C})^{2}$, where $C=\{i,j\}$ or $\{i,j,l\}$ represents the index of pairwise or three-body couplings, $w_{C}$ is a true coupling strength, i.e. $W^{(2)}_{ij}$ or $W^{(3)}_{ijl}$, $\hat{w}_{C}$ is its inferred value, and $M=N(N-1)$ or $N(N-1)(N-2)/2$ is the maximal number of pairwise or three-body couplings, respectively. The mean squared error obtained by the proposed method was lower than that obtained by LASSO and OLS (Table 4). As demonstrated in Figure 4, OLS tends to overfit the coupling strength for uncoupled pairs or triples, which is a possible reason for the large error. This result indicates that the proposed method outperforms LASSO and OLS in inferring the coupling strength.
Figure 4: Inferring coupling strength from time series. Inferred coupling strengths ($\hat{W}^{(2)}_{ij}$ and $\hat{W}^{(3)}_{ijl}$) obtained by the three methods (Left: proposed method, Centre: LASSO, and Right: OLS) are compared with the true strengths ($W^{(2)}_{ij}$ and $W^{(3)}_{ijl}$). The points on the first quadrant represent true positives, while those on the non-zero $x$-axis and $y$-axis represent false negatives and false positives, respectively. The points above (below) the diagonal line indicate an overestimation (underestimation) of the coupling strength. The circle and triangle points represent the inferred results for pairwise and three-body interactions, respectively.
Application to a larger population of oscillators
As demonstrated in previous subsections, the proposed method can accurately identify the interaction type (e.g. pairwise or three-body) and infer the couplings from time series using synthetic data sets with a small number of oscillators (Fig. 2: 12 oscillators and Figs. 3, and 4: 10 oscillators). In order to investigate the feasibility of the proposed method, which involves a larger system in a more realistic scenario, we will simulate a group of 100 oscillators with sparse couplings ($q^{(2)}=0.5$ and $q^{(3)}=0.02$) and log-normally distributed coupling strengths. Figure 5A shows an inference result obtained with the proposed method. The number of errors was low: the number of false positives and false negatives was 8 (${\rm FPR}=0.081$) and 0 (${\rm FNR}=0.000$) for pairwise couplings, and 6 (${\rm FPR}=0.001$) and 0 (${\rm FNR}=0.000$) for three-body couplings, respectively. The MCC was 0.920 and 0.937 for the pairwise and three-body couplings, respectively, indicating that our method can accurately identify the coupling network. Furthermore, the proposed method achieves an accurate inference of the coupling strength (Fig. 5B). The mean squared error $\epsilon^{2}$ of the coupling strength was $5.75\times 10^{-6}$ and $5.76\times 10^{-6}$ for the pairwise and three-body couplings, respectively. These results suggest that the proposed method is applicable to time series obtained from a larger population of oscillators.

### v2

# Distinguishing Pairwise and Higher-Order Interactions in Coupled Oscillators from Time Series

@@ Abstract
Rhythmic phenomena, which are ubiquitous in biological systems, are typically modelled as systems of coupled limit cycle oscillators. Recently, there has been an increased interest in understanding the impact of higher-order interactions on the population dynamics of coupled oscillators. Meanwhile, the estimation of a mathematical model from experimental data is an essential step in understanding the dynamics of real-world complex systems. In coupled oscillator systems, identifying the type of interaction (e.g. pairwise or three-body) is challenging, because different interactions can exhibit similar dynamical states in experimental conditions. In this study, we have developed a method based on the adaptive LASSO (Least Absolute Shrinkage and Selection Operator) to infer the interactions among oscillators from time series data. The proposed method successfully identifies the type of interaction and estimates the probabilities of pairwise and three-body couplings. Through systematic analysis of synthetic datasets, we have demonstrated that our method outperforms two baseline methods, LASSO and OLS (Ordinary Least Squares), in accurately inferring the topology and strength of couplings between oscillators. Furthermore, the proposed method is applied to human brain network data, demonstrating its practical utility. Finally, we extend the method to more general oscillatory systems, including those exhibiting the deformation of limit cycles and those with four-body interactions. These results suggest that our method is a promising tool for identifying interaction mechanisms in oscillatory systems.
@@ Introduction
A complex system can be modelled as a network of interacting nodes or vertices (i.e. elements) connected by links or edges (i.e. individual interaction or couplings). However, in many systems, such as social, ecological, and biological systems, the effects of interaction between groups of three or more elements (higher-order interactions) cannot be ignored. Recently, there has been a significant amount of research on the effects of higher-order interactions between the elements on the network dynamics [5, 9, 10]. In coupled oscillator systems, it has been reported that higher-order interactions can emerge from pairwise interaction, e.g., via the nonlinearity of the coupling [4, 30] and the dependence of the amplitude on the phase [8]. In addition, higher-order interactions can arise spontaneously due to frequency resonance of oscillators [41]. There is also extensive research on the effects of higher-order interactions on dynamics [5]. Higher-order interactions in coupled oscillators exhibit richer dynamical properties compared to pairwise interactions, including multi-stability [56], chaos [7] and explosive synchronisation [51]. Furthermore, higher-order interactions can induce dynamical states such as synchronisation, asynchronisation, and cluster states, whose bifurcation structure is mathematically distinct from pairwise interactions [51, 29, 33].
@@
Estimating a mathematical model from experimental data is an essential step in understanding the dynamics of complex systems in the real world. In the context of coupled oscillator systems, previous studies have developed methods for identifying phase models from data. These methodologies include estimating the phase response curves [18, 60, 37], phase coupling functions [47, 25, 58, 53] (see also review [54, 48]), and asymptotic phases [62, 39, 65, 64]. While several studies have proposed methods for estimating the interaction of coupled oscillator systems [45, 43, 42, 35], these are based on the assumption of pairwise interactions. A key open question is whether it is possible to infer the type of interactions (e.g. pairwise, three-body) in real-world systems. Recently, several methods for estimating higher-order interactions in complex dynamical systems have been proposed [13, 32, 40, 15]. However, identifying the specific type of interaction—whether elements are coupled via pairwise (two-body), three-body, or higher-order group interactions—purely from data remains a significant challenge. In particular, determining whether most pairwise couplings are absent (i.e., effectively zero) based on experimental data is a non-trivial task. In coupled oscillator systems, both higher order interactions [51, 29, 33] and pairwise interactions with additional mechanisms [1] exhibit rich behaviours such as synchronous states, asynchronous states, and cluster states. While the bifurcation structure can differ significantly between pairwise interactions and higher-order interactions, it is not straightforward to determine the interaction type from an experiment. This is because, in experiments, we are unable to fully observe all system variables or to arbitrarily control the bifurcation parameters, and measurements are subject to noise. It is therefore essential to develop a method for identifying the type of interaction and inferring the topology and strength of the couplings between the oscillators.
@@
In this study, we address the following question: "Is it possible to identify the interaction type of coupled oscillators from the data?" We focus on coupled oscillator systems that include pairwise and three-body interactions. We develop a method based on adaptive LASSO [66] for inferring the topology and strength of the individual couplings from time series by assuming the coupling function is known. Most existing network reconstruction methods rely on either OLS (Ordinary Least Squares) [45, 43, 42, 32, 35] or LASSO (Least Absolute Shrinkage and Selection Operator) [20, 50, 59, 15]. In contrast, our proposed method employs the adaptive LASSO, which satisfies desirable oracle properties for large datasets [66]. We then systematically evaluate the performance of the proposed method using synthetic datasets with different interaction types, coupling probabilities, and observation durations, and compare its performance against two baseline methods: OLS and LASSO. Furthermore, we demonstrate the practical applicability of our method by inferring human brain networks [52]. Finally, we present extensions of the method to more general oscillatory systems, including those with limit cycle deformations and those with four-body interactions.
Results
@@
Figure 1: Synchronization dynamics of weakly coupled oscillators in the presence of white noise (Eq. 1). Left: Time series of the Kuramoto order parameter $r$ in a weakly synchronized state for three types of interactions; (a) Pairwise interaction only, (b) Three-body interaction only, and (c) Mixture interactions, where the system has both pairwise and three-body couplings. The average values of the order parameter were $r=0.33,0.25$, and $0.32$ for the pairwise, three-body, and mixture interactions, respectively. Right: Snapshots of the phase distribution of the oscillators when they are synchronized at similar levels ($r=0.67,0.61$ and $0.67$ for the pairwise, three-body, and mixture interaction, respectively). The weights ($K_{2}$, $K_{3}$) and the coupling probabilities ($q^{(2)}$, $q^{(3)}$) were set as follows: A. $K_{2}=0.1$, $q^{(2)}=0.1$and $q^{(3)}=0.0$, B. $K_{3}=0.3$, $q^{(2)}=0.0$, and $q^{(3)}=0.1$, and C. $K_{2}=0.1$, $K_{3}=0.3$, $q^{(2)}=0.05$ and $q^{(3)}=0.05$.
@@
Adaptive LASSO method for inferring the interaction network from time series
@@
[…]
@@
where $h$ is the sampling interval. The minimization is performed by finding the local minimum at which the derivative is zero.
Next, the estimate is refined through minimizing the cost function:
[…]
Figure 2 compares the inferred couplings from the proposed method with those from two baseline methods, LASSO (Least Absolute Shrinkage and Selection Operator) and OLS (Ordinary Least Squares) with a statistical test (see Method for details). LASSO is a popular method for network inference problems in general [20](and others). It has also been applied to inference problems of the same type as ours, namely network inference from time-series [15]. Extensions of LASSO, such as Signal LASSO [50], have been successfully applied to network inference problems. However, to our knowledge, the adaptive LASSO has not been used before. The proposed method was able to infer the topology of the pairwise and three-body couplings with complete accuracy (Fig. 2, Proposed), with no false positives and no false negatives. OLS produced a considerable number of false positives for both pairwise and three-body couplings (Fig. 2, OLS). When the true interaction was pairwise only, the number of false positives was 16 (12 %) and 23 (3.5 %) for pairwise and three-body couplings, respectively. While LASSO gave a lower number of false positives than OLS, it gave the false positives when the true interaction was pairwise (Fig. 2, LASSO). When the true interaction was pairwise only, the number of false positives was 0 (0 %) and 12 (1.8 %) for pairwise and three-body couplings, respectively. Furthermore, the width of the arrow indicates that LASSO underestimates the coupling strength. These results suggest that the proposed method performs better than OLS and LASSO in inferring the couplings from the time series. In the next subsection, we systematically compare the inference performance of the proposed method with that of the baseline methods.
@@
Table 1 shows the accuracy of inferring the interaction type from the time series, which were generated from 20 populations of each pairwise, three-body, and mixture interaction. All the methods achieve an accuracy of over 80 % in identifying the interaction type. In particular, the proposed method and LASSO achieve perfect identification of the interaction type, i.e. 100 % accuracy. Table 2 shows the inferred probabilities of the existence of pairwise and three-body couplings ($\hat{q}^{(2)}$ and $\hat{q}^{(3)}$) obtained from the same data as in Table 1. In the case of synthetic data with pairwise interaction, the proposed method demonstrates the highest accuracy in inferring the coupling probability. In contrast, LASSO underestimated and OLS overestimated the probability of pairwise coupling. In addition, both LASSO and OLS exhibited an overestimation of the probability of three-body coupling by 0.6-0.7 %, and the naive OLS exhibited an overestimation of more than 9.0 %. In the case of synthetic data with three-body interaction, LASSO was the most accurate in inferring the probability of the pairwise couplings, while the proposed method was the most accurate in inferring that of the three-body couplings. In contrast, Naive OLS overestimated the probability of pairwise (5.1 %) and three-body (3.7 %) couplings, and OLS underestimated the probability of the three-body couplings. In the case of synthetic data with mixture interaction, LASSO achieved the most accurate inference for the probability of the pairwise couplings, while the proposed method achieved the most accurate inference for that of the three-body couplings. OLS overestimates the probability of the pairwise couplings, and it underestimates that of the three-body couplings. Naive OLS overestimated the probability of pairwise (7.3 %) and three-body (5.3 %) couplings. Overall, the naive OLS method cannot infer the coupling probabilities in the present situation. In contrast, a recent study [32] has shown that OLS method can infer the higher-order interactions from time series accurately. The discrepancy between these findings can be attributed to the difference in the situation being considered. Previous research has focused on scenarios without noise, while this study considers a system of noisy coupled oscillators. Tables 2 indicates that sophisticated methods, including OLS with a statistical test and LASSO [57] (see Method for details), are essential for distinguishing between weak couplings and no couplings in noisy situations. In the following, the proposed method is compared with two sophisticated methods, i.e. OLS with statistical test and LASSO.
@@
where $N_{\rm TP}$, $N_{\rm TN}$, $N_{\rm FP}$, and $N_{\rm FN}$ are the numbers of true positives (TPs), true negatives (TNs), false positives (FPs), and false negatives (FNs) of the pairwise and three-body couplings, respectively. MCC values range from -1 to 1. If all the couplings are correctly inferred, the MCC value is 1. If the couplings are randomly inferred with the true coupling probability, the expected value of MCC is 0. Figure 3 shows the dependence of the MCC on the observation duration for the three interaction types (pairwise, three-body, and mixture) with three coupling probabilities: $q=0.05,0.10$, and $0.15$. As the coupling probability increases, a longer observation period is required for an accurate inference. For example, 50 cycles are sufficient to accurately ($MCC>0.8$) infer the couplings for the synthetic data of a sparse interaction ($p=0.05$). In contrast, more than 100 cycles are necessary to infer the couplings for the synthetic data of a dense interaction ($p=0.15$). Overall, when the observation period is sufficiently long, the proposed method demonstrates superior performance in terms of accuracy compared to LASSO and OLS. Furthermore, the proposed method outperforms LASSO and OLS in inferring the couplings across a broad range of coupling strengths, $K_{2}$ and $K_{3}$ (see Supplementary Note 1).
@@
Figure 3: Effect of observation time on inference accuracy. Performance of coupling inference (MCC) was calculated for the three types of interaction: (a) Pairwise ($q^{(2)}=q$, $q^{(3)}=0$), (b) Three-body ($q^{(2)}=0$, $q^{(3)}=q$), and (c) Mixture of pairwise and three-body interactions ($q^{(2)}=q/2$, $q^{(3)}=q/2$). The parameter of the coupling probability $q$ was set to 0.05, 0.10 and 0.15, on the left, centre and right panel, respectively.
@@
Finally, we evaluate the inference performance obtained from the coupled oscillators with highly heterogeneous coupling strength. Such interactions are observed in neural circuits in the brain [12, 23]. We simulate a population of oscillators (1) whose coupling strength is given by the Erdös-Rényi random graph with a log-normal distribution of the strength. Figure 4 shows the inference result of the coupling strength by the three methods (our proposed method, LASSO, and OLS). True negatives are plotted at the origin, false negatives are plotted on the positive part of the $x$-axis, false positives are plotted on the positive part of the $y$-axis, and true positives are plotted in the first quadrant. Again, the performance of coupling inference is evaluated using MCC (Table 3). The MCC values of the proposed method were higher than those of LASSO and OLS, indicating that the proposed method outperforms LASSO and OLS in terms of coupling inference. Furthermore, we examine the performance in inferring the coupling strengths. The coupling strengths inferred by OLS are distributed above the diagonal line, indicating that OLS overestimates the strength. Conversely, the coupling strengths inferred by LASSO are distributed below the diagonal line, indicating that LASSO underestimates the strength. The performance in inferring the strength is evaluated by calculating the mean squared error: $\epsilon^{2}=\frac{1}{M}\sum_{C}(w_{C}-\hat{w}_{C})^{2}$, where $C=\{i,j\}$ or $\{i,j,l\}$ represents the index of pairwise or three-body couplings, $w_{C}$ is a true coupling strength, i.e. $W^{(2)}_{ij}$ or $W^{(3)}_{ijl}$, $\hat{w}_{C}$ is its inferred value, and M=N(N1)M=N(N-1) or N(N1)(N2)/2N(N-1)(N-2)/2 is the maximal number of pairwise or three-body couplings, respectively. The mean squared error obtained by the proposed method was lower than that obtained by LASSO and OLS (Table 4). As demonstrated in Figure 4, OLS tends to overfit the coupling strength for uncoupled pairs or triples, which is a possible reason for the large error. This result indicates that the proposed method outperforms LASSO and OLS in inferring the coupling strength.
[…]